solve by the substitution method
6x+5y=1
x=11-3y

Found 2 solutions by nic0le116, oberobic:
Answer by nic0le116(25)

Since you see x=11-3y, you will take what x equals (11-3y) and replace x in the other with this to solve for y.
---------------------------
Take 6x+5y=1
Replace x with (11-3y)
6(11-3y)+5y=1 (distribute 6)
66-18y+5y=1 (combine like terms)
66-13y=1 (get the term attached to y by itself. subtract 66 from each side)
-13y=-65 (get y alone, divide both side by -13)
y=5
------------------------

Now that you have the value of y, you need to replace y in one of the equations to get the value of x.
6x+5(5)=1
6x+25=1
6x=-24
x=-4
------------------------
ANSWERS=
x=-4
y=5

Answer by oberobic(2304)

6x +5y = 1
x = 11 -3y
.
substitute x = 11-3y
.
6(11-3y) +5y = 1
66 -18y +5y = 1
-13y = -65
y = 5
.
substitute y=5 into the first equation
.
6x + 5(5) = 1
6x = -24
x = -4
.
check answer by substituting into the second equation
.
11-3y = 11-15 = -4
x = -4
correct.
.
Answer: x=-4, y=5
.
Done.